trains a and b are traveling in the same direction on parallel tracks. train a is traveling at 60 mph and train b is traveling at 70 mph. train a passes a station at 2:25 pm. if train b passes the same station at 2:37 pm at what time will train b catch up to train a?
Answer by richwmiller(17219)

60*(b+1/5)=d
70b=d
60*(b+1/5)=70b
12=10b
1.2=b
2:37+ 1:12=3:49 pm
check
60*(1.2+.2)=d
70*1.2=84
60*(1.4)=84
ok